OpenStudy (anonymous):

Which equation should be used to solve this problem? Juan has $8 less than Julio. Together they have $48. How much does each have? Let J = Julio's money. Let J - 8 = Juan's money.

OpenStudy (anonymous):

what are your answer choices

OpenStudy (anonymous):

J + J - 8 = 48

OpenStudy (anonymous):

j-(j+8)=48,48-8=j+j,j+(j-8)=48,j(j+8)=48

OpenStudy (anonymous):

i think 48-8=j+j

OpenStudy (anonymous):

thank j+j-8=48 was correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A Chevy is traveling 10 mph faster than a Toyota. The Chevy is traveling at a rate of 60 mph. Let T = the the speed of the Toyota. Which formula below would correctly calculate the speed of the Toyota?

OpenStudy (anonymous):

t=60-10

OpenStudy (anonymous):

thats not one of the answers they are t+10=60,t+(t-10)=60,t+(t+10)=65,t-10=65

OpenStudy (anonymous):

t+10=60

OpenStudy (anonymous):

t + 10 = 60

OpenStudy (anonymous):

THANKS EVERYONE

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:) thats what we are here for